# Language-based Valence and Arousal Expressions between the United States and China: a Cross-Cultural Examination

### v2

@@ Abstract
Although affective expressions of individuals have been extensively studied using social media, research has primarily focused on the Western context. There are substantial differences among cultures that contribute to their affective expressions. This paper examines the differences between Twitter (X) in the United States and Sina Weibo posts in China on two primary dimensions of affect - valence and arousal. We study the difference in the functional relationship between arousal and valence (so-called V-shaped) among individuals in the US and China and explore the associated content differences. Furthermore, we correlate word usage and topics in both platforms to interpret their differences. We observe that for Twitter users, the variation in emotional intensity is less distinct between negative and positive emotions compared to Weibo users, and there is a sharper escalation in arousal corresponding with heightened emotions. From language features, we discover that affective expressions are associated with personal life and feelings on Twitter, while on Weibo such discussions are about socio-political topics in the society. These results suggest a West-East difference in the V-shaped relationship between valence and arousal of affective expressions on social media influenced by content differences. Our findings have implications for applications and theories related to cultural differences in affective expressions.
@@
Young-Min Cho Dandan Pang Stuti Thapa Garrick Sherman Lyle Ungar Louis Tay Sharath Chandra Guntuku University of Pennsylvania Bern University of Applied Sciences University of Tulsa Purdue University {jch0,ungar,sharathg}@seas.upenn.edu, dandan.pang@bfh.ch, stt0912@utulsa.edu garricktsherman@gmail.com, stay@purdue.edu
@@ 1 Introduction
Subjective expression of affect (how we feel) plays a crucial role in understanding learning outcomes in individuals Hourihan et al. (2017), their perceptions Gorn et al. (2001), well-being Xu et al. (2015), and mental and physical health Cohen and Pressman (2006). Multiple theoretical and empirical works have therefore studied the underlying dimensions of affect and their relationship. While there are several models of affective structure, Russell’s 2-dimensional circumplex model is one of the popular models Russell (1980), where orthogonal valence and arousal are considered its most fundamental dimensions Yik et al. (1999). Valence is the measure of pleasure (or displeasure) and arousal of activation or sense of energy.
@@
Understanding the relationship between valence and arousal is interesting from empirical, psychometric, and theoretical perspectives. While previous studies suggest several models to describe the valence arousal relationship Ortony et al. (1990); Lang (1994), where valence and arousal are correspondingly the horizontal and vertical axis in the circumplex, representing the polarity and intensity of affective states, a “V-shaped” relationship of arousal as a function of valence is one of the most widely tested and accepted Kuppens et al. (2013); Cacioppo and Gardner (1999). Arousal is shown to be directly related to the intensity of positive or negative valence with a positivity offset and a negativity bias, with varying levels of cross-cultural support Kuppens et al. (2017a).
While affective structure and relationship are considered universal, most previous works studying the valence arousal relationship are based exclusively on Western samples and do not consider cross-cultural heterogeneity Tsai et al. (2006). Culture can have a big impact on one’s emotional life. Different cultures value emotions (ideal affect) and have different emotional display standards Matsumoto (1990). For example, Americans appreciate enthusiasm (high arousal) to express high positive valence, while Asians more often choose quiet (low arousal) Tsai et al. (2006). Past research also suggests that culture has an important influence in predicting affect perceptions, which can reflect differences in cultural expectations Guntuku et al. (2015b); Zhu et al. (2018); Guntuku et al. (2015a). Furthermore, the few studies that revealed significant evidence of the affective differences across Western and Eastern cultures were based on small samples recruited in lab-based environments Kuppens et al. (2017b). Asking participants to recall recent events and report their emotions can contain measurement bias and other concerns relating to self-report Tarrant et al. (1993); Winograd and Neisser (2006). To counter this, there is a call for more research going beyond self-reports to behaviors Baumeister et al. (2007). The strength of our approach is to examine the generalizability of past cross-cultural findings using actual behaviors of language expressions.
This paper has two contributions. First, we compare the functional relationship between valence and arousal in public affective expression using large-scale social media data, going beyond self-report surveys. Second, we use language samples collected from the United States (in the West) and China (in the East) to elucidate the differences and similarities in the relationship between valence and arousal across cultures by mining ecological expressions from individual timelines on Twitter (X) and Sina Weibo and use thematic content analysis to understand the language markers contributing to such differences and similarities.
Social media data are known to capture individuals’ feelings in a naturalistic and ecological setting and have been shown to reliably estimate well-being Jaidka et al. (2020), sentiment Preoţiuc-Pietro et al. (2016), psychological traits such as personality Schwartz et al. (2013), and health Merchant et al. (2019). The variations in content and user behavior on Weibo and Twitter have been explored in various scenarios Ma (2013); Lin et al. (2016b). Despite the difficulties of working with a non-random, non-representative sample of social media users, posts can reveal a variety of psychological qualities and consequences, including users’ demographics Sap et al. (2014); Zhang et al. (2016), personality Li et al. (2014); Quercia et al. (2011), location Salehi et al. (2017); Zhong et al. (2015), stress levels Guntuku et al. (2019b); Lin et al. (2016a), and mental health Guntuku et al. (2019d); Tian et al. (2018) on both platforms. Examples of posts from the two platforms are shown in Figure 1.
@@ 2 Methods
2.1 Data Collection
Our data consist of public messages posted on Weibo and Twitter. To collect Twitter data, we used the survey platform Qualtrics, which included demographic questions such as gender and age. Participants shared their Twitter handles after completing the survey. Users were compensated for participating, and we received informed consent to access their Twitter posts. There were 3,113 users in the Twitter dataset, with around 3.6 million posts until 2016.
Weibo, unlike Twitter, does not offer an API tool for obtaining random samples over time. So, starting with a random set of individuals from a public dataset Guntuku et al. (2019a), Weibo posts were gathered using a breadth-first search method on users. We obtained over 29 million posts from 2014 from 859,054 people on Weibo. Gender and age were collected from self-reported demographic information on their Weibo profile. Subsetting to users with more than 500 words and with a reasonable age (<100 years) and gender, the dataset consisted of 668,257 Weibo posts from 8,731 users. 500 words were found to be the minimum threshold to obtain reliable psychological estimates from individuals’ language Eichstaedt et al. (2021); Jaidka et al. (2018).
Based on the gender and age distribution of Weibo and Twitter users, we built propensity-score-based matched samples, resulting in 2,191 users each on both platforms with at least 500 words. These matched users had 2.4 million posts on Twitter and 177,042 posts on Weibo. In our matched dataset 67.1% self-reported as being female and 32.9% as male, and the mean age was 26.9 (s.d. 8.8). On Twitter, there were on average 15.6 (s.d. 2.8) words per user and Weibo had 57.3 (s.d. 15.4) words per user.
2.2 Language Preprocessing
@@
2.3 Feature Extraction
We extract two open-vocabulary features from Twitter and Weibo: n-grams and topics. We extract contiguous sequences of one or two words (1-2 grams, Kern et al., 2014; Andrew Schwartz et al., 2013) with pointwise mutual information (PMI = 3; Church and Hanks, 1990). This resulted in unique unigrams and bigrams set of 10,477 for Weibo and 12,798 for Twitter. We extracted the normalized distribution of the n-grams for each post in the Weibo and Twitter datasets. We then used 2000 topics generated using Latent Dirichlet Allocation (LDA, Blei et al., 2003) as the second feature to represent users’ language in our Twitter and Weibo datasets. We utilized topics generated on much larger datasets to favor high diversity and coverage. 2,000 English topics generated a corpus of approximately 18 million Facebook updates with alpha set to 0.30 to favor fewer topics per document. 2,000 Mandarin topics were generated on the entire set of 29 million Weibo posts with similar parameters set in Mallet Andrew Schwartz et al. (2013). Inherently, each topic is realized as a set of words with probabilities. Every post is thus scored in terms of its probability of containing each of the 2000 topics, , which is derived from their probability of containing a word, , and the probability of the words being in given topics, .
2.4 Valence-Arousal Measurement
The circumplex and vector models of emotion have been broadly used for representing affective states Russell (1980); Bradley et al. (1992). In these two-dimensional models, valence is the x-axis, expressing pleasantness and unpleasantness, attractiveness and aversiveness, joy, and sorrow Frijda (1986). Arousal is the y-axis, describing the degree of wakefulness, boredom, excitement, and calm. These models allow any affective state, emotion, word, or expression to be represented as a point in the space, regardless of the difference in language, country, or culture. We measure valence and arousal using a validated data-driven lexicon generated based on the circumplex model in both English and Mandarin. We used NRC Valence, Arousal, and Dominance (NRC-VAD) Lexicon Mohammad (2018) for Twitter data and its translated version for Weibo data. NRC-VAD consists of valence and arousal weights for more than 20,000 words in English and shows a "boomerang" relationship between two dimensions: extremely positive or negative valence is usually paired with high arousal, while calmness matches low arousal. We subtract 0.5 from all scores to make them zero-centered.
Multilinguality is another reason to choose NRC-VAD as our valence-arousal measurement lexicon. There are over 100 languages available for NRC-VAD by translating English terms using Google Translate (August 2022), and the authors claim that most affective norms are stable across languages. Since an original-translated term pair has the same scores, this lexicon avoids the annotator agreement and scale-matching issue, which are common problems using two different lexica over two languages.
2.5 Valence-Arousal Relationship Models
@@
We use Akaike Information Criterion (AIC; Bozdogan, 1987) and Bayesian Information Criterion (BIC; Schwarz, 1978) for model selection. AIC is defined as:
BIC has the following format:
where is the maximized value of the likelihood function of the model, k is the number of parameters, and N is the number of observations. One advantage of using BIC is that it can be used to approximate posterior probability for each model:
While applying the six models, we use mixed effects models to fit the datasets. We assume there is a fixed relationship between valence and arousal across all posts, while the average level of arousal may vary from user to user. The regression models can correctly represent the relationship between the two variables by setting within-person differences as the random effect.
2.6 Differential Language Analysis
We use ordinary least squares (OLS) regression to identify significant associations between the two feature sets, namely n-grams and topics, and valence and arousal. We use feature sets (n-grams or topics) as input, and valence and arousal as outputs and built one OLS model per feature set. We calculated correlation coefficients of each feature dimension and utilized Benjamini-Hochberg p-correction Benjamini and Hochberg (1995) to correct for multiple comparisons and used for indicating meaningful correlations.
@@ 3 Results
3.1 Valence-Arousal Relation Models
| Dataset | Model | AIC | BIC | PostP |
| Model 1 | 0 | |||
| Model 2 | 0 | |||
| Model 3 | 0 | |||
| Model 4 | 0 | |||
| Model 5 | 0 | |||
| Model 6 | 1 | |||
| Model 1 | 0 | |||
| Model 2 | 0 | |||
| Model 3 | 0 | |||
| Model 4 | 0 | |||
| Model 5 | 0 | |||
| Model 6 | 1 |
@@ 3.1 Valence-Arousal Relation Models
Among the different models we tested across Twitter and Weibo data, Model 6 with within-person intercept and slope best fit with the lowest AIC and BIC (Table 1). Within-person models were also significantly different from the models without within-person effects. The presence of an asymmetric V-shape in the data, including a negativity bias and negativity offset, was confirmed in the models on both Twitter and Weibo data. Compared with Weibo, Twitter shows a larger intercept gap (Twitter: ; Weibo: ). The intensity of emotion gets significantly stronger with higher positivity/negativity. This conclusion is consistent with both Twitter and Weibo, with the smallest BIC values in Model 6, characterized by a V shape (Twitter: , Weibo: ) and negativity bias (Twitter: , Weibo: ). The Twitter model has a steeper slope on both positive and negative valence compared to Weibo (Twitter: , ; Weibo: , ).
3.2 Differential Language Analysis
@@
To uncover the content differences in emotional expression across cultures, we utilized differential language analysis to obtain the most correlated n-grams and topics in each platform. Figure 3 shows the top significantly correlated words and phrases with valence and arousal in both platforms. On the dimension of valence, Twitter users tended to use words conveying superlatives (’great’, ’awesome’, ’amazing’) and festive celebrations (’birthday’, ’Christmas’, ’new’, ’win’) in expressing positive valence, while profanity (’shit’, ’fuck’), negation (’hate’, ’bad’, ’wrong’) and discomfort (’wait’, ’tired’, ’stop’) were indicative of negative valence. Conversely, Weibo users commonly employed terms related to personal affect (’like’, ’love’, ’happiness’) and emojis (’oh’, ’heart’) when expressing positive valence, whereas words indicative of negation (’no’) and sorrow (’sad’, ’cry’) are prevalent in expressing negative valence. On the dimension of arousal, Twitter users expressed profanity (’shit’, ’fuck’) and interpersonal expressions (’awesome’, ’amazing’) for high arousal while using terms indicating low activities (’sleep’, ’bed’) and time-oriented description (’today’, ’week’, ’day’, ’time’) for low arousal. In contrast, Weibo users predominantly utilized positive emojis(’steal-laugh’, ’applaud’) to convey high arousal, while employing affirmation (’yes’), negation (’no’), and sharing aspects of daily life (’home’, ’sleep’) to express low arousal.
@@
Comparing the two DLA results for topics, Twitter users had relaxing weekend (’weekend’, ’awesome’, ’amazing’, ’great’, ’retreat’), celebration of events (’birthday’, ’wishes’, ’happy’, ’present’, ’wished’), luck and achievement (’win’, ’won’, ’contest’, ’prize’, ’lottery’) for positive valence high arousal. Conversely, Weibo users discussed affectionate bonding (’love’, ’hopeless’, ’willing’, ’protective’, ’friendly’, where hopeless means love in deep) to express their feelings, particularly in the context of festivals and celebrations (’new year’, ’red envelope’) and interests in celebrities and TV shows (’celebraties’,’singer’). For positive valence low arousal, Twitter users usually talked about relaxing routines (’day’, ’today’, ’good’, ’chilled’) and sleep (’night’, ’sleep’, ’tonight’, ’rest’, ’hoping’). Besides, Weibo users shared family reunion (’home’, ’return’, ’mother’, ’family’, ’new year’) and savory cuisines (’dish’, ’meat’, ’delicious’, ’soup’, ’dish’). When expressing strong negative feelings, Twitter users mainly used profanities (’fucking’, ’fuck’, ’shit’, ’pissed’, ’bullshit’) to convey intense emotions, while Weibo users discussed law enforcement and criminal investigation (’police’, ’crime’, ’suspect’, ’caught’, ’case’). Additionally, Weibo discussions on negative high arousal included the use of emojis (’sweat’) and negative emotions (’schocking’, ’hurt’, ’give up’). Concerning negative valence low arousal, Twitter users usually showed personal negative feelings like tiredness (’tired’, ’sleepy’, ’sleep’, ’sooo’, ’ugh’), engaged in discussions about daily activities (’hair’, ’cut’, ’short’, ’haircut’, ’cutting’) and mentioned words related to time (’hour’, ’minute’). Similarly, Weibo users also mentioned sleep (’sleep’, ’awake’, ’bed’).

### v4

@@ Abstract
While affective expressions on social media have been extensively studied, most research has focused on the Western context. This paper explores cultural differences in affective expressions by comparing valence and arousal on Twitter/X (geolocated to the US) and Sina Weibo (in Mainland China). Using the NRC-VAD lexicon to measure valence and arousal, we identify distinct patterns of emotional expression across both platforms. Our analysis reveals a functional representation between valence and arousal, showing a negative offset in contrast to traditional lab-based findings which suggest a positive offset. Furthermore, we uncover significant cross-cultural differences in arousal, with US users displaying higher emotional intensity than Chinese users, regardless of the valence of the content. Finally, we conduct a comprehensive language analysis correlating n-grams and LDA topics with affective dimensions to deepen our understanding of how language and culture shape emotional expression. These findings contribute to a more nuanced understanding of affective communication across cultural and linguistic contexts on social media.
@@
Young-Min Cho1 Dandan Pang2 Stuti Thapa3 Garrick Sherman1 Lyle Ungar1 Louis Tay4 Sharath Chandra Guntuku1 1University of Pennsylvania 2Bern University of Applied Sciences 3University of Tulsa 4Purdue University {jch0,ungar,sharathg}@seas.upenn.edu, dandan.pang@bfh.ch, stt0912@utulsa.edu garricktsherman@gmail.com, stay@purdue.edu
@@ 1 Introduction
Subjective expressions of affect (how we feel) play a crucial role in understanding learning outcomes in individuals Hourihan et al. (2017), their perceptions Gorn et al. (2001), well-being Xu et al. (2015), and mental and physical health Cohen and Pressman (2006). Multiple theoretical and empirical works have, therefore, examined the underlying dimensions of affect and their relationships. While there are several models of affective structure, Russell’s two-dimensional circumplex model is the most widely recognized, where orthogonal valence (pleasant to unpleasant) and arousal (high to low activation) are represented as the horizontal and vertical axes Russell (1980); Yik et al. (1999).
@@
Understanding the functional relationship between valence and arousal in this two-dimensional space is of empirical, psychometric, and theoretical interest. Among various models from previous studies Ortony et al. (1990); Lang (1994), a "V-shaped" relationship, where arousal is a function of valence, is one of the most widely tested and accepted Kuppens et al. (2013); Cacioppo and Gardner (1999). Arousal is shown to be directly related to the intensity of positive or negative valence with a positivity offset and a negativity bias, with varying levels of cross-cultural support Kuppens et al. (2017a).
While the affective structure and the valence-arousal relationship are often considered universal, most previous studies focus exclusively on Western samples, overlooking cross-cultural heterogeneity Tsai et al. (2006a). Different cultures value emotions (ideal affect) uniquely and adhere to distinct standards for emotional expression Matsumoto (1990). For instance, Americans tend to associate enthusiasm (high arousal) with positive valence, while Asians often prefer quietness (low arousal)Tsai et al. (2006a). Although some studies have highlighted affective differences between Western and Eastern cultures, these are typically based on small, lab-based samplesKuppens et al. (2017b), which may be biased by self-reporting and recall issues Tarrant et al. (1993); Winograd and Neisser (2006).
To address these limitations, researchers advocate for studies that go beyond self-reports and focus on behaviors Baumeister et al. (2007), especially on social media. Social media data offer a naturalistic and ecological setting to capture individuals’ emotions and, despite potential social desirability bias, have been shown to reliably estimate well-being Jaidka et al. (2020); Liou et al. (2023), sentiment Preoţiuc-Pietro et al. (2016), and personality Schwartz et al. (2013); Havaldar et al. (2024).
This paper examines the cross-cultural difference on the affective expressions with functional relationship between valence and arousal by analyzing natural language expressions from Twitter/X (geolocated to the US) and Sina Weibo (in Mainland China) posts. The pipeline of our work is shown in Table 1. The study has three key contributions:
We evaluate functional representations between valence and arousal on large-scale social media data, identifying a negative offset in contrast to previous lab-based studies.
We demonstrate cross-cultural differences in valence and arousal, showing that US users exhibit stronger emotional intensity (higher arousal) than Chinese users across both positive and negative valence.
We employ comprehensive language analysis, correlating n-grams and LDA topics with valence and arousal, providing insights into the functional relationship and cultural divergence in affective expression.
@@ 2 Methods
2.1 Data Preparation
Our data consist of public messages posted on Weibo and Twitter. The content and behavior variations on Weibo and Twitter have been studied in different contexts Ma (2013); Lin et al. (2016). While working with non-random, non-representative samples poses challenges, social media posts can still reveal psychological traits, demographics Sap et al. (2014); Zhang et al. (2016), location Salehi et al. (2017); Zhong et al. (2015), and mental health Guntuku et al. (2019c); Tian et al. (2018).
To collect Twitter data, we used the survey platform Qualtrics, which included demographic questions such as gender and age. Participants from the US shared their Twitter handles after completing the survey. Users were compensated for their time, and consented to share their Twitter posts. There were 3,113 Twitter users, with around 3.6 million posts until 2016.
Weibo, unlike Twitter, does not offer an API tool for obtaining random samples over time. So, starting with a random set of individuals from a public dataset Guntuku et al. (2019a), Weibo posts were gathered using a breadth-first search method on users. We obtained over 29 million posts from 2014 from 859,054 people on Weibo. Gender and age were collected from self-reported demographic information on their Weibo profile. Subsetting to users with more than 500 words and with a reasonable self-reported age (<100 years) and gender, the dataset consisted of 668,257 Weibo posts from 8,731 users. 500 words were found to be the minimum threshold to obtain reliable psychological estimates from individuals’ language Eichstaedt et al. (2021); Jaidka et al. (2018).
Based on the gender and age distribution of Weibo and Twitter users, we built propensity-score-based matched samples, resulting in 2,191 users each on both platforms with at least 500 words. These matched users had 2.4 million posts on Twitter and 177,042 posts on Weibo. In our matched dataset 67.1% self-reported as being female and 32.9% as male, and the mean age was 26.9 (s.d. 8.8). On Twitter, there were on average 15.6 (s.d. 2.8) words per user and Weibo had 57.3 (s.d. 15.4) words per user. The differences in word counts are driven by the lack of character limits to posts on Weibo.
@@ 2.1 Data Preparation
2.2 Valence & Arousal Measurement
The circumplex and vector models of emotion have been broadly used for representing affective states Russell (1980); Bradley et al. (1992). In these two-dimensional models, valence is the x-axis, expressing pleasantness and unpleasantness, attractiveness and aversiveness, joy, and sorrow Frijda (1986). Arousal is the y-axis, describing the degree of wakefulness, boredom, excitement, and calm. These models allow any affective state, emotion, word, or expression to be represented as a point in the space, regardless of the difference in language, country, or culture.
We measure valence and arousal using a validated data-driven lexicon generated based on the circumplex model in both English and Mandarin. We used NRC Valence, Arousal, and Dominance (NRC-VAD) Lexicon Mohammad (2018a) for Twitter data and its translated version for Weibo data. NRC-VAD consists of valence and arousal weights for more than 20,000 words in English and shows a "V-shaped" relationship between two dimensions: extremely positive or negative valence is usually paired with high arousal, while calmness matches low arousal. We subtract 0.5 from all scores to make them zero-centered.
Multilinguality is another reason to choose NRC-VAD as our valence-arousal measurement lexicon. There are over 100 languages available for NRC-VAD (August 2022), and the authors claim that most affective norms are stable across languages. Since an original-translated term pair has the same scores, this lexicon avoids the annotator agreement and scale-matching issue, which are common problems using two different lexica over two languages.
We calculate the valence and arousal scores for each post on Twitter and Weibo using NRC-VAD lexica. For each post, we sum the result of item-wise multiplication of relative word frequency within the post, and the corresponding valence or arousal score for the word. In detail, we follow the formula:
Where represents a post, is a word, is the frequency of word in post , and are the valence scores for post (for annotation) and word (from the NRC-VAD lexicon), respectively, and is the total number of words in post . Arousal is calculated in a similar manner.
2.3 Evaluation of Functional Relationship
@@
We use Akaike Information Criterion (AIC; Bozdogan, 1987), Bayesian Information Criterion (BIC; Schwarz, 1978), posterior probability, and Conditional for model selection. AIC is defined as , and BIC has the following format: , where is the maximized value of the likelihood function of the model, k is the number of parameters, and N is the number of observations. One advantage of using BIC is that it can be used to approximate posterior probability for each model:
While applying the six models, we use mixed effects models to fit the datasets. We assume there is a fixed relationship between valence and arousal across all posts, while the average level of arousal may vary from user to user. The regression models can correctly represent the relationship between the two variables by setting within-person differences as the random effect. In model comparison, to cover both fixed and random effects, we use conditional to represent the proportion of variance explained by the entire model.
2.4 Social Media Language Analysis
Feature Extraction
We extract two open-vocabulary features from Twitter and Weibo: n-grams and topics. We extract contiguous sequences of one or two words (1-2 grams, Kern et al., 2014; Andrew Schwartz et al., 2013) with pointwise mutual information (PMI = 3; Church and Hanks, 1990). This resulted in unique unigrams and bigrams set of 10,477 for Weibo and 12,798 for Twitter. We extracted the normalized distribution of the n-grams for each post in the Weibo and Twitter datasets. We then used 2000 topics generated using Latent Dirichlet Allocation (LDA, Blei et al., 2003) as the second feature to represent users’ language in our Twitter and Weibo datasets. We utilized topics generated on much larger datasets to favor high diversity and coverage. 2,000 English topics generated a corpus of approximately 18 million Facebook updates with alpha set to 0.30 to favor fewer topics per document. These have been shown to perform well across multiple platforms Eichstaedt et al. (2015). 2,000 Mandarin topics were generated on 29 million Weibo posts with similar parameters set in Mallet. Inherently, each topic is realized as a set of words with probabilities. Every post is thus scored in terms of its probability of containing each of the 2000 topics, , which is derived from their probability of containing a word, , and the probability of the words being in given topics, .
Differential Language Analysis
To understand the functional relationship and cultural difference in valence and arousal, we utilized ordinary least squares (OLS) regression to model valence and arousal on post level, controlling for gender and age by matching these variables between the Twitter and Weibo samples. The inputs for the regression were different language features independently extracted from social media posts - n-grams and topics; and outputs were valence and arousal, each of which was constructed using a separate OLS model. We calculated correlation coefficients of each feature dimension and utilized Benjamini-Hochberg p-correction Benjamini and Hochberg (1995) to correct for multiple comparisons and used for indicating meaningful correlations and to control for false discovery rate in multiple hypothesis testing.
The formula for n-gram models are:
where is a post, is an n-gram, is the frequency of in , is coefficients and is a total number of n-grams in the post.
The formula for LDA topic models are:
where is the probability of m belonging to topic t. Models for arousal can be expressed in the same fashion.
@@ 3 Results
3.1 Valence-Arousal Functional Relationship
The full comparison between 6 models are shown in Figure 9. Among the different models we tested across Twitter and Weibo data, Model 6 with within-person intercept and slope best fit with the lowest AIC and BIC, and highest Conditional . Within-person models were also significantly different from the models without within-person effects.
| Dataset | Model | AIC | BIC | PostP | |
| Model 1 | 0 | 0.015 | |||
| Model 2 | 0 | 0.030 | |||
| Model 3 | 0 | 0.050 | |||
| Model 4 | 0 | 0.113 | |||
| Model 5 | 0 | 0.135 | |||
| Model 6 | 1 | 0.139 | |||
| Model 1 | 0 | 0.021 | |||
| Model 2 | 0 | 0.025 | |||
| Model 3 | 0 | 0.030 | |||
| Model 4 | 0 | 0.057 | |||
| Model 5 | 0 | 0.082 | |||
| Model 6 | 1 | 0.084 |
@@ 3.1 Valence-Arousal Functional Relationship
As shown in Figure 2, the presence of an asymmetric V-shape in the data, including a negativity bias and negativity offset, was confirmed in the models on both Twitter and Weibo data. Compared with Weibo, Twitter shows a larger intercept gap (Twitter: ; Weibo: ). The intensity of emotion gets significantly stronger with higher positivity/negativity. This conclusion is consistent with both Twitter and Weibo, with the smallest BIC values in Model 6, characterized by a V shape (Twitter: , Weibo: ) and negativity bias (Twitter: , Weibo: ). The Twitter model has a steeper slope on both positive and negative valence compared to Weibo (Twitter: , ; Weibo: , ).
3.2 Social Media Language Analysis
@@
To uncover the content differences in emotional expression across cultures, we utilized differential language analysis to obtain the most correlated n-grams and topics in each platform. Figure 3 shows the top significantly correlated words and phrases with valence and arousal in both platforms. On the dimension of valence, Twitter users tended to use words conveying superlatives (’great’, ’awesome’, ’amazing’) and festive celebrations (’birthday’, ’Christmas’, ’new’, ’win’) in expressing positive valence, while profanity (’shit’, ’fuck’), negation (’hate’, ’bad’, ’wrong’) and discomfort (’wait’, ’tired’, ’stop’) were indicative of negative valence. Conversely, Weibo users commonly employed terms related to personal affect (’like’, ’love’, ’happiness’) and emojis (’oh’, ’heart’) when expressing positive valence, whereas words indicative of negation (’no’) and sorrow (’sad’, ’cry’) are prevalent in expressing negative valence. On the dimension of arousal, Twitter users expressed profanity (’shit’, ’fuck’) and interpersonal expressions (’awesome’, ’amazing’) for high arousal while using terms indicating low activities (’sleep’, ’bed’) and time-oriented description (’today’, ’week’, ’day’, ’time’) for low arousal. In contrast, Weibo users predominantly utilized positive emojis(’steal-laugh’, ’applaud’) to convey high arousal, while employing affirmation (’yes’), negation (’no’), and sharing aspects of daily life (’home’, ’sleep’) to express low arousal. The version of Figure 3 without using words in NRC-VAD is shown in the Appendix.
@@
We further compare Twitter and Weibo’s DLA results for topics in Figure 4 and 5. Twitter users had relaxing weekend (’weekend’, ’awesome’, ’amazing’, ’great’, ’retreat’), celebration of events (’birthday’, ’wishes’, ’happy’, ’present’, ’wished’), luck and achievement (’win’, ’won’, ’contest’, ’prize’, ’lottery’) for positive valence high arousal. Conversely, Weibo users discussed affectionate bonding (’love’, ’hopeless’, ’willing’, ’protective’, ’friendly’, where hopeless means love in deep) to express their feelings, particularly in the context of festivals and celebrations (’new year’, ’red envelope’) and interests in celebrities and TV shows (’celebrities’,’singer’). For positive valence low arousal, Twitter users usually talked about relaxing routines (’day’, ’today’, ’good’, ’chilled’) and sleep (’night’, ’sleep’, ’tonight’, ’rest’, ’hoping’). Besides, Weibo users shared family reunion (’home’, ’return’, ’mother’, ’family’, ’new year’) and savory cuisines (’dish’, ’meat’, ’delicious’, ’soup’, ’dish’). When expressing strong negative feelings, Twitter users mainly used profanities (’fucking’, ’fuck’, ’shit’, ’pissed’, ’bullshit’) to convey intense emotions, while Weibo users discussed law enforcement and criminal investigation (’police’, ’crime’, ’suspect’, ’caught’, ’case’). Additionally, Weibo discussions on negative high arousal included the use of emojis (’sweat’) and negative emotions (’shocking’, ’hurt’, ’give up’). Concerning negative valence low arousal, Twitter users usually showed personal negative feelings like tiredness (’tired’, ’sleepy’, ’sleep’, ’sooo’, ’ugh’), engaged in discussions about daily activities (’hair’, ’cut’, ’short’, ’haircut’, ’cutting’) and mentioned words related to time (’hour’, ’minute’). Similarly, Weibo users also mentioned sleep (’sleep’, ’awake’, ’bed’).